Soybean meal is 14% protein: cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280 lbs mixture that is 8% protein?
Answer by rfer(16322)

x=soybean=40 lbs
280-x=cornmeal=240 lbs
----------------
.14x+.07(280-x)=.08*280
.14x+19.6-.07x=22.4
.07x=2.8
x=40